OpenStudy (anonymous):

How do I write y = x^2 + 14x + 29 in general form?

Directrix (directrix):

The standardized form of a quadratic equation is ax2 + bx + c = 0, where "a" does not equal 0. [As a review, if the highest degree in an equation is 1, meaning that the x-term is x1 or in the form ax + by = c or y = mx + b, the equation is always linear.] http://jwilson.coe.uga.edu/emt668/EMAT6680.Folders/Barron/unit/Lesson%206/6.html

OpenStudy (anonymous):

lol, my mistake, here. . . http://zonalandeducation.com/mmts/functionInstitute/linearFunctions/lgf.html

Directrix (directrix):

That said, I would go with 1* x ^2 + (14) x + 29 = 0. The queston now is whether standard form and general form mean the same.

OpenStudy (anonymous):

Standard Form and General Form is different.

OpenStudy (anonymous):

I just dont know what general form is.

Directrix (directrix):

What is the difference? Do you have a definition? I'm now thinking that y = 1* x ^2 + (14) x + 29.

OpenStudy (anonymous):

All i know is that its considered rewriting a quadratic equation into general form.

OpenStudy (anonymous):

look to my link Mrbonez( lol. nice nym)

OpenStudy (anonymous):

So general equation is y = (-A/B)x + (-C/B)

Directrix (directrix):

Take a look at this: Find the general form of the family of quadratics with zeroes at x = 1 and x = 3. Subtracting, I get x – 1 = 0 and x – 3 = 0, so the factors were x – 1 and x – 3. However, I can't tell if they divided anything off to get the listed solutions. For instance, they may have started with factors like 4x – 4 or 5x – 15; I can't tell from only the zeroes. The general form of the family of quadratics (that is, the formula for every possible quadratic that has these zeroes) has to include any possible divided-off numbers. For this, I will use a letter to represent the divided-off numbers. The factors I know about are x – 1 and x – 3. Multiplying them together gives me x2 – 4x + 3. Since some number may have been divided out, I have to multiply it back in: The general form is a(x2 – 4x + 3) http://www.purplemath.com/modules/fromzero.htm

OpenStudy (anonymous):

Interesting... thanks

Directrix (directrix):

There seems to be lack of consistency about the distinction between standard and general form. I'm interested now because I don't know. Is there anything in your math text about it? I'll look at my text just in case.